# Multi-Objective Coverage Bayesian Optimization (MOCOBO)

### v2

@@ Abstract
In multi-objective black-box optimization, the goal is typically to find solutions that optimize a set of $T$ black-box objective functions, $f_{1},\ldots f_{T}$, simultaneously. Traditional approaches often seek a single Pareto-optimal set that balances trade-offs among all objectives. In this work, we consider a problem setting that departs from this paradigm: finding a small set of $K<T$ solutions, that collectively “covers" the $T$ objectives. A set of solutions is defined as “covering" if, for each objective $f_{1},\ldots f_{T}$, there is at least one good solution. A motivating example for this problem setting occurs in drug design. For example, we may have $T$ pathogens and aim to identify a set of $K<T$ antibiotics such that at least one antibiotic can be used to treat each pathogen. To address this problem, we propose Multi-Objective Coverage Bayesian Optimization (MOCOBO), a principled algorithm designed to efficiently find a covering set. We validate our approach through experiments on challenging high-dimensional tasks, including applications in peptide and molecular design, where MOCOBO is shown to find high-performing covering sets of solutions. The results show that the coverage of the $K<T$ solutions found by MOCOBO matches or nearly matches the coverage of $T$ solutions obtained by optimizing each objective individually. Furthermore, in in vitro experiments, the peptides found by MOCOBO exhibited high potency against drug-resistant pathogens, further demonstrating the potential of MOCOBO for drug discovery.
@@ 1 Introduction
However, in many real-world scenarios, the trade-offs between objectives can become extreme, with some objectives being nearly mutually exclusive. This makes it virtually impossible to achieve satisfactory performance across all objectives simultaneously. Figure 1 provides a $1D$ example of this demonstrating how the Pareto-optimal points may fail to achieve acceptable performance on all objectives. Such scenarios motivate an interesting generalization of the multi-objective setting that arises in situations where we can propose a small set of solutions that together optimize the multiple objectives and avoid the implicit constraints imposed by extreme trade-offs.
@@ 2 Background and related work
Local Bayesian optimization [17], including TuRBO-$M$, is one recent successful approach to high-dimensional BO. TuRBO-$M$ uses $M$ parallel local optimization runs, each maintaining its own dataset $\mathcal{D}_{i}$ and surrogate model. Each local optimizer proposes candidates only within a hyper-rectangular trust region $\mathcal{T}_{i}$. The trust region $\mathcal{T}_{i}$ is a rectangular area of the input space $\mathcal{X}$, centered on the incumbent $\mathbf{x}^{+}_{i}$. Its side length $\ell_{i}$ is bound by $[\ell_{min},\ell_{max}]$. If an optimizer improves the incumbent for $\rho_{succ}$ consecutive iterations, $\ell_{i}$ expands to $\min(2\ell_{i},\ell_{max})$. If not improved in $\rho_{fail}$ iterations, $\ell_{i}$ is halved. Optimizers are restarted if $\ell_{i}$ drops below $\ell_{min}$. While TuRBO-$M$ is not directly applicable to our problem, we follow methods like MORBO [11] and ROBOT [45] in adapting the use of $M$ coordinated trust regions ($K$ in our case) to a different problem setting.
@@
To our knowledge, the coverage optimization problem setting considered in this paper has not been previously explored in the Bayesian optimization literature. However, similar problem settings have been studied in the context of gradient-based multi-objective optimization. In particular, Ding et al. [13], Li et al. [39], and Lin et al. [40] have investigated related problems in this domain. Ding et al. [13] demonstrate that the coverage optimization problem generalizes various clustering methods, including k-means clustering, and propose an algorithm that extends k-means++ [1] and Lloyd’s algorithm [43] to efficiently solve the coverage problem using gradient descent. Li et al. [39] introduce Many-objective Multi-solution Transport (MosT), a novel framework that scales multi-objective gradient-based optimization to a large number of objectives. MosT uses an optimal transport to establish a balanced assignment between objectives and solutions, allowing each solution to focus on a specific subset of objectives and ensuring that the collective set of solutions covers all objectives. Lin et al. [40] introduce a novel Tchebycheff set (TCH-Set) scalarization approach for covering multiple conflicting objectives using a small set of collaborative solutions. They further propose a smooth Tchebycheff set (STCH-Set) scalarization method to handle non-smoothness in TCH-Set scalarization. While these three papers address related coverage optimization problems, their proposed methods rely on differentiable objective functions, making them unsuitable for the black-box optimization problems considered in this paper. Liu et al. [42] were the first to consider the coverage problem in the setting of black-box optimization. They propose a clustering-based swarm optimization algorithm (CluSO) for the coverage problem in the setting of multi-objective black-box optimization. In Section 4, we provide empirical results demonstrating that our Bayesian optimization approach (MOCOBO) consistently outperforms CluSO by a large margin across all tasks considered.
@@ 3.1 Multi-Objective Coverage Bayesian Optimization (MOCOBO)
In this section, we propose MOCOBO - an algorithm which extends Bayesian optimization to the problem setting above. On each step of optimization $s$, we use our current set of all data evaluated so far $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ to define the best covering set $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\in D_{s}\right\}$ found so far. Here $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$ and $n$ is the number of data points evaluated so far at step $s$. Following Equation 2, we define $S^{*}_{D_{s}}$ as follows:
@@ 3.1.2 Greedy approximation of $S^{*}_{D_{s}}$
The candidate acquisition method in Section 3.1.1 utilizes the best covering set of $K$ points among all data collected so far, $S^{*}_{D_{s}}$. On each step of optimization $s$, MOCOBO must therefore construct $S^{*}_{D_{s}}$ from all observed data $D_{s}$, as $S^{*}_{D_{s}}$ may change on each step of optimization after new data is evaluated and added to $D_{s}$.
@@
Let $T,K$ be finite positive integers such that $K<T$. Let $f_{1},\ldots,f_{T}$ be real valued functions. Let $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ be a dataset of $n$ real valued data points such that for all $\mathbf{x}_{i}$, $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$. Let $S^{*}_{D_{s}}$ be the optimal covering set of size $K$ in $D_{s}$ as defined in Equation 3. Then, constructing $S^{*}_{D_{s}}$ is NP-hard.
@@
See full proof of Lemma 3.1 in Section A.9. ∎
@@
  Compute the initial coverage score for $A$: $c(A)\leftarrow 0$.
@@
        Compute marginal coverage of adding $\mathbf{x}$ to $A$:
@@
Since constructing $S^{*}_{D_{s}}$ is NP-hard, we use an approximate construction of $S^{*}_{D_{s}}$ on each step of optimization $s$. We present Algorithm 1 with $O(nKT)$ running time that provides a $(1-\frac{1}{e})$-approximation of $S^{*}_{D_{s}}$.
@@
The set $A^{*}_{D_{s}}$ output by Algorithm 1 satisfies $c(A^{*}_{D_{s}})\geq\left(1-\frac{1}{e}\right)c(S^{*}_{D_{s}})$.
@@
See full proof of Theorem 3.2 in Section A.10. ∎
Algorithm 1 runtime.
The algorithm iterates $K$ times to construct the covering set. In each iteration, it evaluates at most $n$ candidate points, and for each candidate, it computes the incremental coverage score by iterating over $T$ objectives. The runtime is thus $O(K\cdot n\cdot T)$. For practical applications where we can assume relatively small $K$ and $T$, the runtime is approximately $O(n)$. For an empirical evaluation of the wall-clock runtime, see Section A.2.3.
Corollary 3.3 (Algorithm 1 is the best possible approximation of $S^{*}_{D_{s}}$).
@@ 3.1.3 Extending ECI to the batch acquisition setting (q-ECI)
q-ECI gives the expected improvement in coverage score after simultaneously observing the batch of $q$ points $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{q}\right\}$. However, the resulting Monte Carlo expectation would require $O(q\times m)$ evaluations of Algorithm 1. We therefore adopt a more approximate batching strategy for practical use with large $q$ (see Section A.3 for details).
@@ 3.1.4 MOCOBO trust regions
In order to find a set of $K$ solutions, MOCOBO maintains $K$ simultaneous local optimization runs using $K$ individual trust regions. Each local run $k\in\left\{1,\ldots,K\right\}$ aims to find a single solution $\mathbf{x}^{*}_{k}$, which together form the desired set $S^{*}$. As in the original TuRBO paper, trust regions are rectangular regions of the search space $\mathcal{T}_{k}\subseteq\mathcal{X}$ defined solely by their size and center point. We center the $K$ trust regions on the best covering set of solutions observed so far during optimization. In particular, on each optimization step $s$, we construct an approximation of $S^{*}_{D_{s}}$ using Algorithm 1, and center each trust region $\mathcal{T}_{k}$ on the corresponding point $\mathbf{x}^{*(s)}_{k}$ in $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\right\}$. We then use our proposed ECI (or q-ECI when $q>1$) acquisition function to select $q$ candidates for evaluation from within each trust region. The MC approximation of ECI described in Section 3.1.1 can be straightforwardly applied to select candidates in trust region $\mathcal{T}_{k}$ by sampling the $m$ discrete points from within the rectangular bounds of $\mathcal{T}_{k}$. Since we select and evaluate $q$ candidates from each of the $K$ trust regions, the total number of observed data points $n$ at step $s$ of MOCOBO is $n=s\times K\times q$. As in the original TuRBO algorithm, each trust region $\mathcal{T}_{k}$ has success and failure counters that dictate the size of the trust region. For MOCOBO, we count a success for trust region $\mathcal{T}_{k}$ whenever $\mathcal{T}_{k}$ proposes a candidate on step $s$ that improves upon the best coverage score and is included in $S^{*}_{D_{s+1}}$.
@@ 4 Experiments
We evaluate MOCOBO on four high-dimensional, multi-objective BO tasks for which finding a set of $K<T$ solutions to cover the $T$ objectives is desirable. Detailed descriptions of each task are in Section 4.1. Two tasks involve continuous search spaces, allowing direct application of MOCOBO, while the other two involve structured spaces (molecules and peptides), requiring an extension for structured optimization.
@@
We implement MOCOBO using BoTorch [2] and GPyTorch [20]. Code to reproduce all results in the paper is available on GitHub: placeholder. We use an acquisition batch size of $20$ for all tasks and across all BO methods compared. Further implementation details are provided in subsection A.7. For our structured optimization tasks, we apply MOCOBO in the latent space of a pre-trained VAE on which we perform regular end-to-end updates with the surrogate model during optimization [44].
@@
In Figure 2, we plot the best coverage score $c$ Equation 1 obtained by the $K$ best covering solutions found so far after a certain number of function evaluations on each task. Since BO baseline methods are not designed to optimize coverage directly, instead aiming to find a single solution or a set of solutions to optimize the $T$ objectives, we plot the best coverage score $c$ obtained by the $K$ best covering solutions found by the method. All plots show mean coverage scores averaged over $20$ replications of each method, and show standard errors.
@@
Table 1: Best set of $K=4$ peptides found by one run of MOCOBO for the “template free" peptide design task. We provide the APEX model’s predicted MIC for each sequence on each of the $11$ target pathogenic bacteria. The target pathogenic bacteria B1,,$\ldots$,, … , B11 are listed in Table A.7. (-) and (+) indicate Gram negative and Gram positive pathogenic bacteria respectively. The best/lowest MIC achieved for each pathogenic bacteria is in bold in each column. See row “TF1" in Figure 3 left, and Figure A.1 for in vitro MICs for this set of $K=4$ peptides. See Table A.1 for an analogous result on the “template constrained" peptide design task.
@@
As TuRBO, LOL-BO, and ROBOT target single-objective optimization, we conduct $T$ independent runs to optimize each of the $T$ objectives for the multi-objective task. We use all solutions gathered from the $T$ runs to compute the best covering set of $K$ solutions found by each method. Unlike LOL-BO and TuRBO which seek a single best solution for a given objective, a single run of ROBOT seeks a set of $M$ solutions that are pairwise diverse. We run ROBOT with $M=K$ so that each independent run of ROBOT seeks $M=K$ diverse solutions. The aggregate result of the $T$ independent runs for ROBOT is thus $K$ diverse solutions for each of the $T$ objectives. We compare to the best covering $K$ solutions from among those $K*T$ solutions. See Section A.7.4 for more details on the diversity constraints used by ROBOT and the associated hyperparameters. MORBO is a multi-objective optimization method and can thus be applied directly to each multi-objective optimization task. For each run of MORBO, we compare to the best covering set of $K$ solutions found among all solutions proposed by the run. For CluSO, no extension is needed as this method is designed for coverage optimization. We run CluSO with the same $K$ used by MOCOBO, and all other hyperparameters set to CluSO defaults.

### v3

@@ Abstract
In multi-objective black-box optimization, the goal is typically to find solutions that optimize a set of $T$ black-box objective functions, $f_{1},\ldots f_{T}$, simultaneously. Traditional approaches often seek a single Pareto-optimal set that balances trade-offs among all objectives. In this work, we consider a problem setting that departs from this paradigm: finding a small set of $K<T$ solutions, that collectively “covers" the $T$ objectives. A set of solutions is defined as “covering" if, for each objective $f_{1},\ldots f_{T}$, there is at least one good solution. A motivating example for this problem setting occurs in drug design. For example, we may have $T$ pathogens and aim to identify a set of $K<T$ antibiotics such that at least one antibiotic can be used to treat each pathogen. To address this problem, we propose Multi-Objective Coverage Bayesian Optimization (MOCOBO), a principled algorithm designed to efficiently find a covering set. We validate our approach through experiments on challenging high-dimensional tasks, including applications in peptide and molecular design, where MOCOBO is shown to find high-performing covering sets of solutions. The results show that the coverage of the $K<T$ solutions found by MOCOBO matches or nearly matches the coverage of $T$ solutions obtained by optimizing each objective individually. Furthermore, in in vitro experiments, the peptides found by MOCOBO exhibited high potency against drug-resistant pathogens, further demonstrating the potential of MOCOBO for drug discovery. We make code available here: https://github.com/nataliemaus/mocobo.
@@ 1 Introduction
However, in many real-world scenarios, the trade-offs between objectives can become extreme, with some objectives being nearly mutually exclusive. This makes it virtually impossible to achieve satisfactory performance across all objectives simultaneously. Figure˜1 provides a $1D$ example of this demonstrating how the Pareto-optimal points may fail to achieve acceptable performance on all objectives. Such scenarios motivate an interesting generalization of the multi-objective setting that arises in situations where we can propose a small set of solutions that together optimize the multiple objectives and avoid the implicit constraints imposed by extreme trade-offs.
@@ 2 Background and related work
Local Bayesian optimization [17], including TuRBO-$M$, is one recent successful approach to high-dimensional BO. TuRBO-$M$ uses $M$ parallel local optimization runs, each maintaining its own dataset $\mathcal{D}_{i}$ and surrogate model. Each local optimizer proposes candidates only within a hyper-rectangular trust region $\mathcal{T}_{i}$. The trust region $\mathcal{T}_{i}$ is a rectangular area of the input space $\mathcal{X}$, centered on the incumbent $\mathbf{x}^{+}_{i}$. Its side length $\ell_{i}$ is bound by $\ell_{min},\ell_{max}]$. If an optimizer improves the incumbent for $\rho_{succ}$ consecutive iterations, $\ell_{i}$ expands to $\min(2\ell_{i},\ell_{max})$. If not improved in $\rho_{fail}$ iterations, $\ell_{i}$ is halved. Optimizers are restarted if $\ell_{i}$ drops below $\ell_{min}$. While TuRBO-$M$ is not directly applicable to our problem, we follow methods like MORBO [11] and ROBOT [45] in adapting the use of $M$ coordinated trust regions ($K$ in our case) to a different problem setting.
@@
To our knowledge, the coverage optimization problem setting considered in this paper has not been previously explored in the Bayesian optimization literature. However, similar problem settings have been studied in the context of gradient-based multi-objective optimization. In particular, Ding et al. [13], Li et al. [39], and Lin et al. [40] have investigated related problems in this domain. Ding et al. [13] demonstrate that the coverage optimization problem generalizes various clustering methods, including k-means clustering, and propose an algorithm that extends k-means++ [1] and Lloyd’s algorithm [43] to efficiently solve the coverage problem using gradient descent. Li et al. [39] introduce Many-objective Multi-solution Transport (MosT), a novel framework that scales multi-objective gradient-based optimization to a large number of objectives. MosT uses an optimal transport to establish a balanced assignment between objectives and solutions, allowing each solution to focus on a specific subset of objectives and ensuring that the collective set of solutions covers all objectives. Lin et al. [40] introduce a novel Tchebycheff set (TCH-Set) scalarization approach for covering multiple conflicting objectives using a small set of collaborative solutions. They further propose a smooth Tchebycheff set (STCH-Set) scalarization method to handle non-smoothness in TCH-Set scalarization. While these three papers address related coverage optimization problems, their proposed methods rely on differentiable objective functions, making them unsuitable for the black-box optimization problems considered in this paper. Liu et al. [42] were the first to consider the coverage problem in the setting of black-box optimization. They propose a clustering-based swarm optimization algorithm (CluSO) for the coverage problem in the setting of multi-objective black-box optimization. In Section˜4, we provide empirical results demonstrating that our Bayesian optimization approach (MOCOBO) consistently outperforms CluSO by a large margin across all tasks considered.
@@ 3.1 Multi-Objective Coverage Bayesian Optimization (MOCOBO)
In this section, we propose MOCOBO - an algorithm which extends Bayesian optimization to the problem setting above. On each step of optimization $s$, we use our current set of all data evaluated so far $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ to define the best covering set $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\in D_{s}\right\}$ found so far. Here $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$ and $n$ is the number of data points evaluated so far at step $s$. Following Equation˜2, we define $S^{*}_{D_{s}}$ as follows:
@@ 3.1.2 Greedy approximation of $S^{*}_{D_{s}}$
The candidate acquisition method in Section˜3.1.1 utilizes the best covering set of $K$ points among all data collected so far, $S^{*}_{D_{s}}$. On each step of optimization $s$, MOCOBO must therefore construct $S^{*}_{D_{s}}$ from all observed data $D_{s}$, as $S^{*}_{D_{s}}$ may change on each step of optimization after new data is evaluated and added to $D_{s}$.
@@
Let $T,K$ be finite positive integers such that $K<T$. Let $f_{1},\ldots,f_{T}$ be real valued functions. Let $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ be a dataset of $n$ real valued data points such that for all $\mathbf{x}_{i}$, $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$. Let $S^{*}_{D_{s}}$ be the optimal covering set of size $K$ in $D_{s}$ as defined in Equation˜3. Then, constructing $S^{*}_{D_{s}}$ is NP-hard.
@@
See full proof of Lemma˜3.1 in Section˜A.9. ∎
@@
 Compute the initial coverage score for AA: $c(A)\leftarrow 0$.
@@
   Compute marginal coverage of adding $\mathbf{x}$ to AA:
@@
Since constructing $S^{*}_{D_{s}}$ is NP-hard, we use an approximate construction of $S^{*}_{D_{s}}$ on each step of optimization $s$. We present Algorithm˜1 with $O(nKT)$ running time that provides a $(1-\frac{1}{e})$-approximation of $S^{*}_{D_{s}}$.
@@
The set $A^{*}_{D_{s}}$ output by Algorithm˜1 satisfies $c(A^{*}_{D_{s}})\geq\left(1-\frac{1}{e}\right)c(S^{*}_{D_{s}})$.
@@
See full proof of Theorem˜3.2 in Section˜A.10. ∎
Algorithm˜1 runtime.
The algorithm iterates $K$ times to construct the covering set. In each iteration, it evaluates at most $n$ candidate points, and for each candidate, it computes the incremental coverage score by iterating over $T$ objectives. The runtime is thus $O(K\cdot n\cdot T)$. For practical applications where we can assume relatively small $K$ and $T$, the runtime is approximately $O(n)$. For an empirical evaluation of the wall-clock runtime, see Section˜A.2.3.
Corollary 3.3 (Algorithm˜1 is the best possible approximation of $S^{*}_{D_{s}}$).
@@ 3.1.3 Extending ECI to the batch acquisition setting (q-ECI)
q-ECI gives the expected improvement in coverage score after simultaneously observing the batch of $q$ points $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{q}\right\}$. However, the resulting Monte Carlo expectation would require $O(q\times m)$ evaluations of Algorithm˜1. We therefore adopt a more approximate batching strategy for practical use with large $q$ (see Section˜A.3 for details).
@@ 3.1.4 MOCOBO trust regions
In order to find a set of $K$ solutions, MOCOBO maintains $K$ simultaneous local optimization runs using $K$ individual trust regions. Each local run $k\in\left\{1,\ldots,K\right\}$ aims to find a single solution $\mathbf{x}^{*}_{k}$, which together form the desired set $S^{*}$. As in the original TuRBO paper, trust regions are rectangular regions of the search space $\mathcal{T}_{k}\subseteq\mathcal{X}$ defined solely by their size and center point. We center the $K$ trust regions on the best covering set of solutions observed so far during optimization. In particular, on each optimization step $s$, we construct an approximation of $S^{*}_{D_{s}}$ using Algorithm˜1, and center each trust region $\mathcal{T}_{k}$ on the corresponding point $\mathbf{x}^{*(s)}_{k}$ in $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\right\}$. We then use our proposed ECI (or q-ECI when $q>1$) acquisition function to select $q$ candidates for evaluation from within each trust region. The MC approximation of ECI described in Section˜3.1.1 can be straightforwardly applied to select candidates in trust region $\mathcal{T}_{k}$ by sampling the $m$ discrete points from within the rectangular bounds of $\mathcal{T}_{k}$. Since we select and evaluate $q$ candidates from each of the $K$ trust regions, the total number of observed data points $n$ at step $s$ of MOCOBO is $n=s\times K\times q$. As in the original TuRBO algorithm, each trust region $\mathcal{T}_{k}$ has success and failure counters that dictate the size of the trust region. For MOCOBO, we count a success for trust region $\mathcal{T}_{k}$ whenever $\mathcal{T}_{k}$ proposes a candidate on step $s$ that improves upon the best coverage score and is included in $S^{*}_{D_{s+1}}$.
@@ 4 Experiments
We evaluate MOCOBO on four high-dimensional, multi-objective BO tasks for which finding a set of $K<T$ solutions to cover the $T$ objectives is desirable. Detailed descriptions of each task are in Section˜4.1. Two tasks involve continuous search spaces, allowing direct application of MOCOBO, while the other two involve structured spaces (molecules and peptides), requiring an extension for structured optimization.
@@
We implement MOCOBO using BoTorch [2] and GPyTorch [20]. Code to reproduce MOCOBO results on all tasks considered in the paper is available on GitHub: https://github.com/nataliemaus/mocobo. We use an acquisition batch size of $20$ for all tasks and across all BO methods compared. Further implementation details are provided in subsection A.7. For our structured optimization tasks, we apply MOCOBO in the latent space of a pre-trained VAE on which we perform regular end-to-end updates with the surrogate model during optimization [44].
@@
In Figure˜2, we plot the best coverage score $c$ Equation˜1 obtained by the $K$ best covering solutions found so far after a certain number of function evaluations on each task. Since BO baseline methods are not designed to optimize coverage directly, instead aiming to find a single solution or a set of solutions to optimize the $T$ objectives, we plot the best coverage score $c$ obtained by the $K$ best covering solutions found by the method. All plots show mean coverage scores averaged over $20$ replications of each method, and show standard errors.
@@
Table 1: Best set of $K=4$ peptides found by one run of MOCOBO for the “template free" peptide design task. We provide the APEX model’s predicted MIC for each sequence on each of the $11$ target pathogenic bacteria. The target pathogenic bacteria $\ldots, B1$1 are listed in Table˜A.7. (-) and (+) indicate Gram negative and Gram positive pathogenic bacteria respectively. The best/lowest MIC achieved for each pathogenic bacteria is in bold in each column. See row “TF1" in Figure˜3 left, and Figure˜A.1 for in vitro MICs for this set of $K=4$ peptides. See Table˜A.1 for an analogous result on the “template constrained" peptide design task.
@@
As TuRBO, LOL-BO, and ROBOT target single-objective optimization, we conduct $T$ independent runs to optimize each of the $T$ objectives for the multi-objective task. We use all solutions gathered from the $T$ runs to compute the best covering set of $K$ solutions found by each method. Unlike LOL-BO and TuRBO which seek a single best solution for a given objective, a single run of ROBOT seeks a set of $M$ solutions that are pairwise diverse. We run ROBOT with $M=K$ so that each independent run of ROBOT seeks $M=K$ diverse solutions. The aggregate result of the $T$ independent runs for ROBOT is thus $K$ diverse solutions for each of the $T$ objectives. We compare to the best covering $K$ solutions from among those KTK*T solutions. See Section˜A.7.4 for more details on the diversity constraints used by ROBOT and the associated hyperparameters. MORBO is a multi-objective optimization method and can thus be applied directly to each multi-objective optimization task. For each run of MORBO, we compare to the best covering set of $K$ solutions found among all solutions proposed by the run. For CluSO, no extension is needed as this method is designed for coverage optimization. We run CluSO with the same $K$ used by MOCOBO, and all other hyperparameters set to CluSO defaults.